# Beyond Wide-Angle Images: Unsupervised Video Portrait Correction via Spatiotemporal Diffusion Adaptation

### v1

@@ Abstract
Wide-angle cameras, despite their popularity for content creation, suffer from distortion-induced facial stretching—especially at the edge of the lens—which degrades visual appeal. To address this issue, we propose an image portrait correction framework using diffusion models named ImagePD. It integrates the long-range awareness of transformer and multi-step denoising of diffusion models into a unified framework, achieving global structural robustness and local detail refinement. Besides, considering the high cost of obtaining video labels, we then repurpose ImagePD for unlabeled wide-angle videos (termed VideoPD), by spatiotemporal diffusion adaption with spatial consistency and temporal smoothness constraints. For the former, we encourage the denoised image to approximate pseudo labels following the wide-angle distortion distribution pattern, while for the latter, we derive rectification trajectories with backward optical flows and smooth them. Compared with ImagePD, VideoPD maintains high-quality facial corrections in space and mitigates the potential temporal shakes sequentially. Finally, to establish an evaluation benchmark and train the framework, we establish a video portrait dataset with a large diversity in people number, lighting conditions, and background. Experiments demonstrate that the proposed methods outperform existing solutions quantitatively and qualitatively, contributing to high-fidelity wide-angle videos with stable and natural portraits. The codes and dataset will be available.
@@ 1 Introduction
Traditional wide-angle portrait correction methods typically require precise camera parameters (e.g., focal length [29]) as a prerequisite, followed by stereographic projection, face and line detection, and energy optimization to achieve geometric correction [29]. Such a pipeline is complex in process and inapplicable when the relevant camera parameters are unknown. The same goes for video portrait correction work [17]. In contrast, learning-based solutions eliminate this issue [33] by directly learning the spatial mapping from wide-angle images to rectification flows in a supervised manner. However, these methods often exhibit noticeable temporal shakes when applied to videos. Moreover, there is currently a lack of datasets for video portrait correction due to the significantly higher cost of labeling video data.
To this end, we take a pioneering step in achieving video portrait correction with diffusion models. First of all, we present a global-to-local portrait correction diffusion model, named ImagePD, to generate high-quality rectification flows, which integrates the advantages of both Transformer and diffusion models. The Transformer model establishes long-range geometric dependencies by capturing spatial global features across entire images, providing structural guidance for the subsequent diffusion process. The diffusion model refines high-fidelity flow patterns through iterative denoising guided by Transformer-derived structural information. This co-design enables structure-to-detail pixel-wise correction, where structural priors from the Transformer and detail-generation capacity from the diffusion model are synergistically combined.
More importantly, we propose an unsupervised spatiotemporal diffusion adaption approach to enable ImagePD the capability to rectify video sequences stably. The repurposed model, which we term VideoPD, addresses the issue of temporal shakes without requiring video correction labels. Concretely, we first use the pre-trained ImagePD to generate rectification flows for each video frame as pseudo-labels. When ImagePD is adapted to unseen wide-angle videos, these pseudo-labels can guide the denoising process, ensuring spatial consistency in facial corrections with the wide-angle distortion distribution prior. To enforce temporal smoothness, we derive rectification trajectories across sequential frames using backward optical flows and design a temporal smoothness constraint to mitigate sequential jitters. By unsupervised spatiotemporal adaptation with diffusion models, VideoPD achieves stable wide-angle video correction that not only preserves high-fidelity facial details but also frees from labor-intensive video annotations.
@@
  • We propose the first portrait correction diffusion model, named ImagePD, which integrates long-range dependences of transformer and iterative refinements of diffusion models for global-to-local rectification.
  • We design an unsupervised spatiotemporal diffusion adaption framework to transfer ImagePD to VideoPD without labeled videos. It combines spatio-temporal optical flows to track correction trajectories, establishing spatial consistency and temporal smoothness constraints.
@@ 2.1 Portrait Correction
Portrait correction in wide-angle images is a significant research area. Traditional methods, such as Shih et al.[29]’s content-aware warping, adjust distortions by analyzing image content and camera parameters. However, these methods bbnnv nbrely on camera parameters or complex inputs and are difficult to extend to video processing. Stereographic projection methods[29] can preserve local conformality but struggle with temporal consistency in videos.
Deep learning has introduced calibration-free solutions. Tan et al.[33] proposed a two-stage neural network leveraging wide-angle images and correction flow datasets. In contrast, Zhu et al. [47] developed a semi-supervised Transformer to reduce annotation costs. Nie et al.’s method CoupledTPS[23], iteratively couples limited TPS models with a warping flow to precisely correct facial features.
Portrait correction in wide-angle videos has garnered less research focus than image-based studies. Current deep learning methods, designed solely for static image processing, do not accommodate the temporal dynamics inherent in video data. This stems from the high cost of manually annotating wide-angle videos, which demands frame-by-frame effort and proves impractical for continuous video tasks, resulting in a scarcity of comprehensive training datasets.
Lai et al.[17]. introduced temporal consistency energy minimization based on traditional spherical projection and energy minimization to optimize videos. However, this method lacks a connection between the global information of wide-angle videos and the detailed information needed for facial correction. It also lacks adaptability and stability when dealing with complex environments and conditions.
To address this issue, we propose learning video correction principles from image networks.
@@ 2.2 Video stabilization
Traditional video stabilization techniques primarily stabilize the video by smoothing the feature trajectories between multiple frames or adjacent frames [35, 18, 28]. Subsequently, grid-based methods [35, 20, 22] and optical flow (Liu et al.[21]; Geo et al.[15]) have also been applied to represent motion, becoming key tools for stabilizing videos. With the rapid development of deep learning, deep learning methods have been applied to video stabilization technology [45, 43, 46, 25, 39], where the original video frames are input, and stabilized video frames are directly output. For example, Wang et al. [34] proposed an end-to-end learning framework, and Xu et al. [37] used adversarial networks to generate target images.DUT [38], which learns video stabilization by simply watching unstable videos. Additional methods, including unsupervised online video stitching (Nie et al.[24]), and meta-learning-based approaches to improve full-frame stabilization (Ali et al.[1]), have been explored to enhance stability. These approaches, ranging from optical flow to deep learning models, provide valuable solutions for addressing video temporal shake in video portrait correction.
@@ 2.3 Diffusion Models
Diffusion models (DMs)[40] have recently achieved remarkable success in generative tasks, extending from image generation to video-related applications. In video generation, early advancements include Video Diffusion Models (VDM)[13], which adapted the 2D U-Net architecture from image DMs to a 3D U-Net for generating coherent video frames. To enhance temporal resolution, Ho et al.[12] introduced a cascade of video super-resolution models, improving the quality of generated sequences. Additionally, methods like Make-A-Video [30] achieved text-to-video generation by leveraging pre-trained text-to-image DMs, avoiding the need for paired text-video data.
More recent work focuses on video editing and enhancing temporal consistency. For example, Tune-A-Video [36] fine-tunes models on single input videos to produce coherent video outputs, while FateZero [26] ensures temporal consistency using attention-based blending techniques. Similarly, Rerender-A-Video [41] and VideoControlNet[14] introduce pixel-aware latent fusion and motion-aware adjustments to maintain visual consistency during edits.
Figure 2: An overview of the proposed framework for portrait distortion correction, integrating Transformer networks and diffusion models. The framework consists of two stages: ImagePD for generating high-quality rectification flow and VideoPD for applying this correction to videos, ensuring spatiotemporal consistency and smooth stabilization.
@@ 3.1 Overview
The overview of our method is shown in Fig. 2, where our framework operates in two cascaded stages: correction and stabilization.In the first stage, ImagePD tackles the distorted image $I_{s}$, producing precise optical flows using a Transformer-guided diffusion approach. The refined optical flow is then applied to generate a geometrically aligned image $\hat{I_{s}}$.
In the second stage, temporal pixel trajectories $T$ are extracted from shaky video sequences, represented as the unstable video frames $V_{s}$. Using the corrected frames $\hat{I}_{s}$ as the reference, the final corrected and stabilized output $V_{\text{final}}$ is computed as:
$V_{\text{final}}=C(V_{s}\mid\hat{I_{s}},T),$ (1)
where $C$ denotes the spatiotemporal fusion module.
3.2 ImagePD
@@
As illustrated in Fig. 2(a,b), our ImagePD is a framework for high-quality portrait correction by predicting precise optical flow fields to rectify facial distortions.
3.2.1 Model Architecture
@@
In the first stage of ImagePD, we focus on capturing the global geometric structure of the distorted input image $I_{s}$. The Transformer network [47], denoted as MS-Unet in Fig. 2(a), with its advantage in capturing long-range information, extracts a set of intermediate layer features $h$ that effectively reflect the global geometric structure of the distorted input image. These features $h$ serve as a structural prior, guiding the diffusion model in the next stage to accurately align the facial geometry while preserving spatial coherence.
Building on the global structure from the first stage, this stage, as shown in Fig. 2(b), fines local details using a diffusion-based approach.
@@
In order to ensure that the correction process primarily focuses on the area of the face, we use InsightFace [27, 2, 3, 6, 5, 7, 8, 10, 11, 9] as a pre-trained model to process the original facial image $I_{\text{s}}$, allowing accurate face detection and keypoint localization. Based on these detection results, a binary mask M is generated, which clearly delineates the facial region. This mask provides spatial guidance for subsequent processing and improves the quality of the correction. Leveraging the detail-handling capability of diffusion models, we form the conditioning input for ImagePD by concatenating $M$, $I_{s}$, and $h$:
$x_{in}=cat[M,I_{s},h]$ (2)
Our approach leverages the iterative refinement capabilities of Denoising Diffusion Implicit Models (DDIM)[31], enabling iterative data generation through noise-adding and denoising processes. In DDIM, an original sample $x_{0}$ is gradually corrupted with Gaussian noise over $T$ timesteps, eventually approximating a standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$. The reverse process starts from $x_{T}$ and iteratively removes noise to recover a clean sample, modeled as a learned denoising transition.
A time-conditioned U-Net, denoted as $\epsilon_{\theta}$, is trained to predict the noise $\epsilon$ added at each step. During training, noise is added to $x_{0}$ to generate intermediate states $x_{t}$, and the model optimizes the following mean squared error (MSE) loss:
$L=\mathbb{E}_{t\sim U(1,T),\,x_{0}\sim q(x_{0}),\,\epsilon\sim\mathcal{N}(0,I)}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\Big{]}$ (3)
@@
Adapting this concept, ImagePD redefines the diffusion process to operate directly on optical flow fields.
The task of the diffusion network at this stage is not to directly generate the corrected facial image, but to predict an optical flow field $F^{corr}$. Compared to directly generating an image, the optical flow field offers sub-pixel precision, enabling more fine-grained guidance and correction of minor deformations in the image.
By iteratively refining the predicted optical flow using the diffusion model, our method ultimately produces a highly accurate optical flow prediction. This prediction is then used to warp distortion correction on the facial image, producing the corrected output:
$\hat{I_{s}}=\mathcal{W}(F^{corr},I_{s})$ (4)
@@ 3.2.2 Optimization Objective
$\mathcal{L}_{image}=\lambda_{1}\mathcal{L}_{mask}+\lambda_{2}\mathcal{L}_{photo}+\lambda_{3}\mathcal{L}_{flow}$ (5)
@@
Training objective: $\mathcal{L}_{image}$ comprises three components, each addressing a specific aspect of the correction process.
Optical Flow Loss $\mathcal{L}_{flow}$: Measures the difference between the model-predicted optical flow and the ground truth optical flow.
$\mathcal{L}_{flow}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}W\left(F^{Corr}(h,w)-F_{\text{gt}}(h,w)\right)^{2}$ (6)
Image Loss $\mathcal{L}_{photo}$: Evaluates the difference between the corrected image after optical flow correction and the ground truth image.
$\mathcal{L}_{photo}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}W\left(\hat{I}_{s}(h,w)-I_{\text{gt}}(h,w)\right)^{2}$ (7)
Mask-based Sobel Loss $\mathcal{L}_{mask}$: Uses the Sobel operator to extract edge information and calculates the edge loss within the facial region mask.
@@
$\quad+\left|G_{y}(F^{Corr})-G_{y}(F_{\text{gt}})\right|\big{]}\cdot M$
3.3 VideoPD
Videos are dynamic sequences composed of continuous multi-frame images =$\mathcal{I}=\left\{\mathcal{I}_{1},\mathcal{I}_{2},\dots,\mathcal{I}_{n}\right\}$. To achieve unsupervised wide-angle video correction, we use pre-trained ImagePD to generate corrected optical flow for each frame as pseudo-labels. However, directly applying these pseudo-labels to the video frame sequence for correction results in severe temporal shakes sequentially, significantly decreasing the video’s quality and visual experience. To address this issue, we propose further fine-tuning of ImagePD to enhance its adaptability in distortion video scenarios by introducing VideoPD.
3.3.1 Correction Trajectory
Our method aims to modulate and control the motion between corresponding frames, so we first use the RAFT [44] network to estimate the inter-frame optical flow from the original wide-angle video. For each pair of consecutive frames $I_{t}$ and $I_{t+1}$, RAFT computes the backward optical flow $f_{t\to t+1}$to avoid generating invalid regions. Next, each frame is processed by the ImagePD module, which leverages a diffusion model to predict the optical flow of correction $F^{corr}_{t}$, effectively serving as a pseudo-label. This flow contains detailed adjustment information for the facial regions, addressing distortions introduced by the wide-angle lens.
To address the temporal shake caused by independent corrections between frames, we construct a series of two-dimensional meshes for each frame on the original frames. The pseudo-label optical flow $F^{corr}_{t}$ is then used to adjust the positions of the control points in the mesh and we can obtain an updated mesh $G$.
To ensure smooth corrections across time, the corrected mesh $G$ is used for reprojection onto the optical flow $f_{t\to t+1}$ from RAFT and the pseudo-label flow $F^{corr}_{t}$. This fusion yields a refined temporal optical flow $F^{final}_{t}$, which effectively mitigates warping shakes and ensures consistency between frames. A visual representation of this relationship Eq. 9 is provided in Fig. 3.
$r(t+1)=f_{t\to t+1}\Bigl{(}G+F_{t}^{\mathrm{Corr}}\Bigr{)}+F_{t}^{\mathrm{Corr}}-F_{t+1}^{\mathrm{Corr}}$ (9)

### v2

# Beyond Wide-Angle Images: Structure-to-Detail Video Portrait Correction via Unsupervised Spatiotemporal Adaptation

@@ Abstract
Wide-angle cameras, despite their popularity for content creation, suffer from distortion-induced facial stretching—especially at the edge of the lens—which degrades visual appeal. To address this issue, we propose a structure-to-detail portrait correction model named ImagePC. It integrates the long-range awareness of the transformer and multi-step denoising of diffusion models into a unified framework, achieving global structural robustness and local detail refinement. Besides, considering the high cost of obtaining video labels, we then repurpose ImagePC for unlabeled wide-angle videos (termed VideoPC), by spatiotemporal diffusion adaption with spatial consistency and temporal smoothness constraints. For the former, we encourage the denoised image to approximate pseudo labels following the wide-angle distortion distribution pattern, while for the latter, we derive rectification trajectories with backward optical flows and smooth them. footnotetext: *Corresponding author; † project lead Compared with ImagePC, VideoPC maintains high-quality facial corrections in space and mitigates the potential temporal shakes sequentially in blind scenarios. Finally, to establish an evaluation benchmark and train the framework, we establish a video portrait dataset with a large diversity in the number of people, lighting conditions, and background. Experiments demonstrate that the proposed methods outperform existing solutions quantitatively and qualitatively, contributing to high-fidelity wide-angle videos with stable and natural portraits. The codes and dataset will be available. The video is available at https://wenbo-nie.github.io/structure-to-detail-portrait-correction/.
@@ 1 Introduction
Traditional wide-angle portrait correction methods typically require precise camera parameters (e.g., focal length (Shih, Lai, and Liang 2019)) as a prerequisite, followed by stereographic projection, face and line detection, and energy optimization to achieve geometric correction (Shih, Lai, and Liang 2019). Such a pipeline is complex in process and inapplicable when the relevant camera parameters are unknown. The same goes for video portrait correction work (Lai et al. 2022). In contrast, learning-based solutions eliminate this issue (Tan et al. 2021) by directly learning the spatial mapping from wide-angle images to rectification flows in a supervised manner. However, these methods often exhibit noticeable temporal shakes when applied to videos. Moreover, there is currently a lack of datasets for video portrait correction due to the significantly higher cost of labeling video data.
To this end, we take a pioneering step in achieving video portrait correction without video annotations. First of all, we present a structure-to-detail portrait correction model, named ImagePC, to generate high-quality rectification flows, which integrates the advantages of both Transformer and diffusion models. The Transformer model establishes long-range geometric dependencies by capturing spatial global features across entire images, providing structural guidance for the subsequent diffusion process. The diffusion model refines high-fidelity flow patterns through iterative denoising guided by Transformer-derived structural information. This hybrid architecture enables structure-to-detail pixel-wise correction, where structural priors from the Transformer and detail-generation capacity from the diffusion model are synergistically combined.
More importantly, we propose an unsupervised spatiotemporal adaption approach to enable ImagePC the capability to rectify video sequences stably. The repurposed model, which we term VideoPC, addresses the issue of temporal shakes without requiring video correction labels. Concretely, we first use the pre-trained ImagePC to generate rectification flows for each video frame as pseudo-labels. When ImagePC is adapted to unseen wide-angle videos, these pseudo-labels can guide the denoising process, ensuring spatial consistency in facial corrections with the wide-angle distortion distribution prior. To enforce temporal smoothness, we derive rectification trajectories across sequential frames using backward optical flows and design a temporal smoothness constraint to mitigate sequential jitters. With these spatiotemporal constraints, VideoPC achieves stable wide-angle video correction in blind scenarios that not only preserves high-fidelity facial details but also frees from labor-intensive video annotations.
@@
  • We propose a structure-to-detail portrait correction model, named ImagePC, which integrates long-range dependences of transformer and iterative refinements of diffusion models for global-to-local rectification.
  • We design an unsupervised spatiotemporal adaptation framework to transfer ImagePC to VideoPC without labeled videos. It combines spatio-temporal optical flows to track correction trajectories, establishing spatial consistency and temporal smoothness constraints.
@@ 2.1 Portrait Correction
Portrait correction in wide-angle images is a significant research area. Traditional methods, such as Shih et al.’s(Shih, Lai, and Liang 2019) content-aware warping, adjust distortions by analyzing image content and camera parameters. However, these methods often rely on camera parameters or complex inputs and are difficult to extend to video processing. Stereographic projection methods(Shih, Lai, and Liang 2019) can preserve local conformality but struggle with temporal consistency in videos.
Deep learning has introduced calibration-free solutions. Tan et al.(Tan et al. 2021) proposed a two-stage neural network leveraging wide-angle images and correction flow datasets. In contrast, Zhu et al. (Zhu et al. 2022) developed a semi-supervised Transformer to reduce annotation costs. Nie et al.’s method CoupledTPS(Nie et al. 2024a), iteratively couples limited TPS models with a warping flow to precisely correct facial features.
Existing deep learning methods, designed for static images, fail to address the temporal dynamics of videos due to expensive frame-by-frame annotation and limited datasets. Lai et al.(Lai et al. 2022) proposed a temporal consistency optimization method using spherical projection and energy minimization, but it struggles to integrate global video information with facial correction details and lacks adaptability in complex environments. To address this issue, we propose learning video correction principles from image networks.
@@ 2.2 Video stabilization
Traditional video stabilization techniques primarily stabilize the video by smoothing the feature trajectories between multiple frames or adjacent frames (Wang et al. 2013; Lee et al. 2009; Rublee et al. 2011). Subsequently, grid-based methods (Wang et al. 2013; Liu et al. 2013, 2016; Yan et al. 2024) and optical flow  (Liu et al. 2014);  (James, Jain, and Rajwade 2022) have also been applied to represent motion, becoming key tools for stabilizing videos. With the rapid development of deep learning, deep learning methods have been applied to video stabilization technology (Zhao and Ling 2020; Yu and Ramamoorthi 2019; Zhao et al. 2023; Peng et al. 2024; Yang et al. 2006), where the original video frames are input, and stabilized video frames are directly output. For example, Wang et al.  (Wang et al. 2019) proposed an end-to-end learning framework, and  (Xu et al. 2018) used adversarial networks to generate target images. DUT  (Xu et al. 2022), which learns video stabilization by simply watching unstable videos. Additional methods, including unsupervised online video stitching  (Nie et al. 2024b), and meta-learning-based approaches to improve full-frame stabilization  (Ali et al. 2024), have been explored to enhance stability. These approaches, ranging from optical flow to deep learning models, provide valuable solutions for addressing video temporal shake in video portrait correction.
@@ 2.3 Diffusion Models
Diffusion models  (Yang et al. 2024) excel in generative tasks, extending from image generation to video-related applications. In video generation, early advancements include Video Diffusion Models (VDM) (Ho et al. 2022b), which adapted the 2D U-Net architecture from image DMs to a 3D U-Net for generating coherent video frames. To enhance temporal resolution, Ho et al.(Ho et al. 2022a) introduced a cascade of video super-resolution models, improving the quality of generated sequences. Additionally, methods like Make-A-Video (Singer et al. 2022) achieved text-to-video generation by leveraging pre-trained text-to-image DMs, avoiding the need for paired text-video data.
More recent work focuses on video editing and enhancing temporal consistency. For example, Tune-A-Video (Wu et al. 2023) fine-tunes models on videos for coherent outputs, while FateZero (Qi et al. 2023) ensures temporal consistency using attention-based blending techniques. Similarly, Rerender-A-Video (Yang et al. 2023) and VideoControlNet(Hu and Xu 2023) introduce pixel-aware latent fusion and motion-aware adjustments to maintain visual consistency during editing.
Figure 2: The proposed framework consists of two models. (1) ImagePC: We design a structure-to-detail architecture to generate high-quality rectification flows for each individual frame. (2) VideoPC: We adapt ImagePC from the image to video domain with spatial consistency and temporal smoothness constraints.
@@ 3.1 Overview
The overview of our framework is shown in Fig. 2, consisting of an image correction model and a video correction model. In the image model, ImagePC tackles the distorted image $I_{s}$, producing precise optical flows using a structure-to-detail rectification architecture. The predicted flow is then applied to generate a geometrically corrected image $\hat{I_{s}}$.
In the video model, directly adopting the image model to correct the video frames can easily yield temporal shakes due to temporally unsmooth warps. But we could take the correction flows from the image model as the reference and transfer the image model to the video domain through spatiotemporal adaptation.
3.2 ImagePC
@@
As illustrated in Fig. 2(a,b), our ImagePC is a framework for high-quality portrait correction by predicting precise optical flow fields to rectify facial distortions.
Structure-to-Detail Model
@@
In the first stage of ImagePC, we focus on capturing the global geometric structure of the distorted input image $I_{s}$. The Transformer network (Zhu et al. 2022), in Fig. 2(a), with its advantage in capturing long-range information, extracts a set of intermediate layer features $h$ that effectively reflect the global geometric structure of the distorted input image. These features $h$ serve as a structural prior, guiding the diffusion model in the next stage to accurately align the facial geometry while preserving spatial coherence.
@@
Building on the global structure from the first stage, the next stage, as shown in Fig. 2(b), refines local details using a diffusion-based approach. To ensure that the correction process primarily focuses on the area of the face, we use InsightFace (Ren et al. 2023; Guo et al. 2021; Guo, Zhao, and Wang 2024) to process the original facial image $I_{\text{s}}$, allowing accurate face detection and keypoint localization. Based on these detection results, a binary mask M is generated, which clearly delineates the facial region to enable better portrait perception. This mask provides spatial guidance for subsequent processing and improves the quality of the correction. Leveraging the detail-handling capability of diffusion models, we form the conditioning input for ImagePC by concatenating $M$, $I_{s}$, and $h$:
$x_{in}=cat[M,I_{s},h].$ (1)
Our approach leverages the iterative refinement capabilities of Denoising Diffusion Implicit Models (DDIM)(Song, Meng, and Ermon 2022), enabling iterative data generation through noise-adding and denoising processes.
@@
Adopting this concept, ImagePC redefines the diffusion process to operate directly on optical flow fields.
Unlike traditional image-to-image methods that directly generate a corrected facial image, we modify the diffusion network in this stage to predict an optical flow field $F^{corr}$. The repurposed image-to-flow pipeline not only ensures that the corrected content totally loyal to the original content, but also enables the derivation of video correction trajectories.
With high-quality correction flows, the wide-angle rectification can be formulated as:
$\hat{I_{s}}=\mathcal{W}(F^{corr},I_{s}).$ (2)
@@
$\mathcal{L}_{image}=\lambda_{1}\mathcal{L}_{mask}+\lambda_{2}\mathcal{L}_{photo}+\lambda_{3}\mathcal{L}_{flow},$ (3)
@@
Optical Flow Loss $\mathcal{L}_{flow}$: It directly measures the motion-level difference between the predicted optical flow and the ground truth optical flow as:
$\mathcal{L}_{flow}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}W\left(F^{Corr}(h,w)-F_{\text{gt}}(h,w)\right)^{2}.$ (4)
Image Loss $\mathcal{L}_{photo}$: It evaluates the image-level difference between the corrected image and the ground truth image as:
$\mathcal{L}_{photo}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}W\left(\hat{I}_{s}(h,w)-I_{\text{gt}}(h,w)\right)^{2}.$ (5)
Mask-based Sobel Loss $\mathcal{L}_{mask}$: We use the Sobel kernels ($G_{x}$ and $G_{y}$) to extract edge information and calculate the edge loss within the facial region mask as:
@@
$\quad+\left|G_{y}(F^{Corr})-G_{y}(F_{\text{gt}})\right|\big{]}\cdot M.$
3.3 VideoPC
Videos are dynamic sequences composed of continuous multi-frame images $\mathcal{I}=\left\{\mathcal{I}_{1},\mathcal{I}_{2},\dots,\mathcal{I}_{n}\right\}$. To achieve unsupervised wide-angle video correction, we use pre-trained ImagePC to generate corrected optical flow for each frame as pseudo-labels. However, directly applying these pseudo-labels to the video sequences will result in severe temporal shakes, significantly decreasing the video’s quality and visual experience. To address this issue, we further propose to transfer ImagePC to VideoPC with the adaptability in wide-angle video scenarios.
Correction Trajectory
Our method aims to control the comprehensive motion between consecutive frames. To represent such motions, we derive correction trajectories that are built on inter-frame and intra-frame optical flows. For inter-frame motions, we use the RAFT (Zhang et al. 2024) to capture the temporal flows ($f_{t\to t+1}$ and $f_{t+1\to t}$) from wide-angle video frames ($I_{t}$ and $I_{t+1}$). For intra-frame motions, we leverage the VideoPC model (the same architecture as ImagePC) to predict the rectification flows ($F^{corr}_{t}$). The remaining key question is how to combine these flows into our desired correction trajectories.
Based on Forward Optical Flow
For forward optical flow (i.e., $F^{corr}_{t}$ denotes the correspondences from wide-angle images to rectified images), we can identify the differences of corresponding points within the consecutive rectified results as the following formula:
$r(t+1)=f_{t+1\to t}+F_{t}^{\mathrm{Corr}}\Bigl{(}G+f_{t+1\to t}\Bigr{)}-F_{t+1}^{\mathrm{Corr}},$ (7)
where $G\in\mathbb{R}^{2\times H\times W}$ denotes pixel-wise grid coordinates. The latter part is the position in the (t+1)-th frame (i.e., $F_{t+1}^{\mathrm{Corr}}+G$), while the upper is the corresponding position in the t-th frame (i.e., $f_{t+1\to t}+F_{t}^{\mathrm{Corr}}(G+f_{t+1\to t})+G$).
Based on Backward Optical Flow
However, in the actual portrait distortion correction process, employing forward optical flow to warp is prone to generating invalid spatial holes. Therefore, all current optical flow-based correction methods employ backward optical flows, that is, $F^{corr}_{t}$ denotes the mapping flow from rectified results to original wide-angle images. The difference in optical flow direction makes the relative motion described in Eq. 7 unsuitable for practical backward optical flow transformations. To compare the rectified results of the $t$-th and $(t+1)$-th frames, we need to identify the differences in corresponding positions in the input image for the same grid location. However, directly comparing the two optical flows $F^{corr}_{t}$ and $F^{corr}_{t+1}$ is not feasible, as the two backward optical flows do not correspond to the same points. Therefore, we need to leverage $f_{t\to t+1}$ to correct the temporal misalignment.
The position difference with backward flows is derived as follows:
$r(t+1)=f_{t\to t+1}\Bigl{(}G+F_{t}^{\mathrm{Corr}}\Bigr{)}+F_{t}^{\mathrm{Corr}}-F_{t+1}^{\mathrm{Corr}},$ (8)
where Fig. 3 provide a visual explanation about Eq. 8.